You are considering a bond with a face value of $1 000 and a coupon rate of 2.0%. The bond has 16 year until maturity and coupon payments are paid semiannually. The yield to maturity on similar securities in the market is 8.3% What is the current price of this bond?
Answer #1

Solution:         

Statement showing calculation of Current price of the bond

Sl.No.

Particulars

Period

Cash Flow

(1)

Annuity Factor @ 4.15 %

(2)

Discounted Cash Flow

(3) = (1) * (2)

1

Half yearly Interest

( $ 1,000 * 2 % * (6/12))

1 – 32

$ 10

17.537121

$ 175.37121

2

Maturity Amount

32

$ 1,000

0.272209

$ 272.209

3

Current price of the bond

$ 447.58021

4

Current price of the bond ( when rounded off to two decimal places)

$ 447.58            

Note :

1.Since Interest is payable half yearly and the no. of years to maturity is 16 years, the price per bond is calculated by converting 16 years into (16 *2) = 32 half yearly periods

2.Thus, the Interest earned per period = $ 1000 * 2 % * (6/12) = $ 10

3. Since the Interest is paid semi annually the discount rate used is = 8.3 % * (6/12) = 4.15 %

4. Interest earned during the 32 periods is discounted using PVIFA ( 4.15 % , 32 ) = 17.537121

5.The Present value of $ 1,000 recoverable at maturity is to be calculated using the half yearly discount rate of (8.3 * (6 /12) ) = 4.15 %

6. Thus PVF ( 4.15 % , 32 ) = 0.272209